THE PERIMETER OF A BASKETBALL COURT IS 78 METERS AND THE LENGTH IS 6 METERS LONGER THAN TWICE THE WIDTH. WHAT ARE THE LENGTH AND WIDTH?
Answer by rfer(16322)

let x=width=11
let 2x+6=length=28
2(2x+6)+2x=78
4x+12+2x=78
6x=66
x=11